f the wheel on a car accelerates from rest at a constant rate of 1.5 rev/s2, how fast will it be turning after 3 s?

Select one:

a. 1.5 rev/s

b. 2 rev/s

c. 3 rev/s

d. 4.5 rev/s
Answer #1

Using:

\alpha =\frac{\Delta\omega }{\Delta t}

Where

\alpha is angular acceleration

\Delta \omega = Change in angular velocity

\Delta t = Change in time

Substituing values

1.5 = (v - 0) / 3

Where v is final angualr velocity after 3s

v = 4.5 rev/s

Hence answer is d